OpenStudy (anonymous):

Choose 2 axioms that allows 22 + (m + 8) to be written as m + 30 commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (anonymous):

a+(b+c) = (a+b)+c a+b+c = a+(b+c)